# The Effect of Language Diversity When Fine-Tuning Large Language Models for Translation

### v1

@@ Abstract
Prior research diverges on language diversity in LLM fine-tuning: Some studies report benefits while others find no advantages. Through controlled fine-tuning experiments across 132 translation directions, we systematically resolve these disparities. We find that expanding language diversity during fine-tuning improves translation quality for both unsupervised and—surprisingly—supervised pairs, despite less diverse models being fine-tuned exclusively on these supervised pairs. However, benefits plateau or decrease beyond a certain diversity threshold. We show that increased language diversity creates more language-agnostic representations. These representational adaptations help explain the improved performance in models fine-tuned with greater diversity. [^1] We will release our code and models upon acceptance.
@@ 1 Introduction
General-purpose LLMs like Llama 3 (Grattafiori et al., 2024) show promise for machine translation but require targeted fine-tuning beyond their incidental bilingualism (Briakou et al., 2023) to match the performance of specialized translation systems. Through fine-tuning approaches ranging from two-stage methods (Li et al., 2024; Zeng et al., 2024; Stap et al., 2024) to more sophisticated optimization techniques (Xu et al., 2025; Zhu et al., 2024b), LLMs such as Tower (Alves et al., 2024) now outperform traditional NMT systems (Kocmi et al., 2024; Deutsch et al., 2025).
@@ 2 Fine-tuning and evaluation design
We compare the following incremental fine-tuning approaches using the Tower family of models, which are built on Llama 2 and underwent continued pre-training with a mixture of monolingual and parallel data: base: TowerBase-7B model without task-specific fine-tuning, serving as our baseline.
fsec: base fine-tuned only on fully supervised English-centric translation directions (10 directions), representing minimal supervision.
@@ 3 Results
Expanding from fs+ps+un (132 directions) to 272 directions reveals nuances in the diversity-performance relationship. Unsupervised directions benefit from increased diversity, while fully supervised directions show slight performance decline, suggesting benefits plateau beyond a certain threshold (details in Appendix C.1). This contradicts prior work that found monotonic improvements with diversity Wang et al. (2022); Dang et al. (2024), but aligns with Muennighoff et al. (2023)’s observation of diminishing returns when scaling multilingual pretraining beyond certain language counts.
@@
Figure 2: SVCCA similarity scores between fine-tuned and base models across layers. Lower values indicate greater adaptation during fine-tuning. base-fsec (blue), base-fs (red), and base-fs+ps+un (green) are compared, with their mean shown in black. Shaded regions represent confidence intervals. Middle layers show most significant adaptation, with lowest mean similarity (0.91) at layer 12. fp+ps+un exhibits greater adaptation throughout the network.
@@
Figure 3: Kernel density estimation of layer 12 activations for base (top) and fs+ps+un (bottom). Colors represent translation directions. Intra-cluster distances show increased specialization for single-target clusters in fs+ps+un, while multi-target cluster C1 demonstrates increased overlap.
@@ B.1 Optimization
[…]
@@ Appendix C Additional results
C.1 Scaling diversity to 272 languages
To investigate whether further increasing language diversity yields additional benefits, we compared our most diverse model from the main experiments (fs+ps+un with 132 directions) to an even more diverse setup including 272 translation directions. While maintaining a similar distribution of language families as in our main experiments, we added five additional languages:
  • Germanic family: Danish (da, North Germanic) and Afrikaans (af, West Germanic)
  • Slavic family: Slovak (sk, West Slavic) and Bulgarian (bg, South Slavic)
  • East Asian languages: Vietnamese (vi, different writing system but shares vocabulary with Chinese)
This selection maintains balanced representation across language families while introducing controlled diversity within each family. All additional languages are represented in both NTREX and FLORES-200.
Importantly, we evaluate both models on the same set of languages and directions as used throughout the paper. The additional languages are only used during fine-tuning to increase diversity, allowing us to measure their impact on the original set of translation directions.
Figure 4: COMET-strict scores comparing models trained on 132 directions and 272 directions. Both are evaluated on the original test set with the same language pairs as used throughout the paper. Unsupervised directions show clearest benefits from increased diversity (+0.01), while fully supervised directions show a slight decrease (-0.003), suggesting potential diversity trade-offs.
Figures 4 shows the performance comparison between our 132-direction model (fs+ps+un) and the expanded 272-direction model.
For fully supervised pairs, we observe a slight performance decrease (-0.003 COMET-strict) when scaling to 272 directions. Partially supervised directions show almost identical performance (+0.001), while unsupervised directions demonstrate the clearest benefit (+0.01) from increased language diversity.
These results suggest that language diversity benefits may plateau or even slightly decline for already well-represented language pairs. The slight reduction in fully supervised performance may indicate a trade-off between focused optimization and broader generalization, where extremely high diversity can dilute the model’s effectiveness for specific well-represented languages. Nonetheless, the continued improvements for unsupervised directions (with respect to the original TowerBase model) demonstrate that higher diversity provides additional benefits for these previously unseen language combinations, even though both the 132 and 272 direction models include these pairs during fine-tuning.
C.2 Regularization alone insufficient
@@
Figure 5 shows the COMET-strict scores comparing fs models with weight decay values of 0.01 (standard) and 0.10.555The results for weight decay at 0.05 were very similar to 0.10 and are omitted for clarity. Increasing the regularization strength has minimal impact on translation performance across all language categories. For fully supervised directions, both models achieved identical mean scores (0.875). For partially supervised directions, the difference was negligible (0.699 vs. 0.697). For unsupervised directions, the model with stronger regularization actually performed slightly worse (0.483 vs. 0.490).
@@
C.3 Results not due to multi-parallel data
@@
C.4 Invariance to model scale

### v2

@@ Abstract
Prior research diverges on language diversity in LLM fine-tuning: Some studies report benefits while others find no advantages. Through controlled fine-tuning experiments across 132 translation directions, we systematically resolve these disparities. We find that expanding language diversity during fine-tuning improves translation quality for both unsupervised and—surprisingly—supervised pairs, despite less diverse models being fine-tuned exclusively on these supervised pairs. However, benefits plateau or decrease beyond a certain diversity threshold. We show that increased language diversity creates more language-agnostic representations. These representational adaptations help explain the improved performance in models fine-tuned with greater diversity.
@@ 1 Introduction
General-purpose LLMs like Llama 3 (Grattafiori et al., 2024) show promise for machine translation but require targeted fine-tuning beyond their incidental bilingualism (Briakou et al., 2023) to match the performance of specialized translation systems. Through fine-tuning approaches ranging from two-stage methods (Li et al., 2024; Zeng et al., 2024) to more sophisticated optimization techniques (Xu et al., 2025; Zhu et al., 2024b), LLMs such as Tower (Alves et al., 2024) now outperform traditional NMT systems (Kocmi et al., 2024; Deutsch et al., 2025).
@@ 2 Fine-tuning and evaluation design
We compare the following incremental fine-tuning approaches using the Tower family of models, which are built on Llama 2 and underwent continued pre-training with a mixture of monolingual and parallel data:
  • base: TowerBase-7B model without task-specific fine-tuning, serving as our baseline. fsec: base fine-tuned only on fully supervised English-centric translation directions (10 directions), representing minimal supervision.
@@ 3 Results
Figure 2: COMET-strict scores comparing models trained on 132 directions and 272 directions. Both are evaluated on the original test set with the same language pairs as used throughout the paper. Unsupervised directions show the clearest benefits from increased diversity (+0.01), while fully supervised directions show a slight decrease (-0.003), suggesting potential diversity trade-offs.
@@
To probe the limits of linguistic breadth, we expanded extscfs+ps+un from 132 to 272 directions by adding Danish, Afrikaans, Slovak, Bulgarian, and Vietnamese while preserving the family balance of our original design. These languages appear only during fine-tuning—the evaluation set remains identical to isolate the effect of additional training diversity. Figure 2 contrasts the two setups. Unsupervised directions gain the most (+0.01 COMET-strict), partially supervised pairs remain stable (+0.001), and fully supervised pairs dip slightly (-0.003), revealing a plateau for well-supported languages. This pattern refines earlier reports of monotonic gains with diversity (Wang et al., 2022; Dang et al., 2024) and echoes the diminishing returns seen in multilingual pretraining (Muennighoff et al., 2023).
@@
Findings transfer across architectures
We further replicate our setups on the multilingual Gemma 2 2B model (Team et al., 2025) to test whether diversity benefits extend beyond the Tower family. Table 1 shows the same monotonic improvements we observe for Tower: expanding the fine-tuning directions from fsec to fs to fs+ps+un consistently raises COMET-strict scores across supervision regimes. The smaller architecture starts from stronger zero-shot performance than TowerBase (0.469 vs. 0.253 on unsupervised pairs) yet still gains the most from the full 132-direction setup (fs+ps+un), confirming that language-diversity gains are not model-family artifacts.
Model FS PS UN
base 0.620 0.522 0.469
fsec 0.723 0.618 0.498
fs 0.725 0.645 0.587
fs+ps+un 0.739 0.711 0.697
Table 1: COMET-strict scores for fine-tuning Gemma 2 2B across the supervision regimes. Columns FS, PS, and UN correspond to fully supervised, partially supervised, and unsupervised pairs. Language-diversity progression mirrors the trends observed for Tower, demonstrating cross-architecture robustness.
@@
Figure 3: SVCCA similarity scores between fine-tuned and base models across layers. Lower values indicate greater adaptation during fine-tuning. base-fsec (blue), base-fs (red), and base-fs+ps+un (green) are compared, with their mean shown in black. Shaded regions represent confidence intervals. Middle layers show most significant adaptation, with lowest mean similarity (0.91) at layer 12. fs+ps+un exhibits greater adaptation throughout the network.
@@
[…]
@@ 4 Conclusion
Acknowledgments
This research was funded in part by the Netherlands Organization for Scientific Research (NWO) under project number VI.C.192.080. We thank SURF (www.surf.nl) for the support in using the National Supercomputer Snellius.
@@ B.1 Optimization
We conducted hyperparameter tuning on our development set (FLORES-200 dev), exploring learning rate scheduler $\in\{\text{cosine, inverse square root}\}$, batch size $\in\{128,256\}$, and learning rate $\in\{2\times 10^{-5},2\times 10^{-6}\}$.
@@ Appendix C Additional results
C.1 Regularization alone insufficient
@@
Figure 5 shows the COMET-strict scores comparing fs models with weight decay values of 0.01 (standard) and 0.10.444The results for weight decay at 0.05 were very similar to 0.10 and are omitted for clarity. Increasing the regularization strength has minimal impact on translation performance across all language categories. For fully supervised directions, both models achieved identical mean scores (0.875). For partially supervised directions, the difference was negligible (0.699 vs. 0.697). For unsupervised directions, the model with stronger regularization actually performed slightly worse (0.483 vs. 0.490).
@@
C.2 Results not due to multi-parallel data
@@
C.3 Invariance to model scale